Draw the line/s for each item using

graphing paper

1. Find the distance from the point (2,3) to the line 3x + 4y + 9 = 0.

2. Find the distance between the given line 4x – 3y = 12 and 4x – 3y + 8 = 0. lines.


1
Expert's answer
2021-10-18T03:45:19-0400

1. Distance from point (2,3) to the line 3x+4y+9=0

=>r=*3(2)+4(3)+9))/√3^2+4^2

=>r=(16+12+9)/√9+16

=>r=27/5

=>r= 5.4


2. The distance between two parallel lines, given by equations
    ax + by + c = 0

    ax + by + d = 0

=(-12-8)/(√4^2+-3^2
=20/5
=4 units